OpenStudy (anonymous):

Look at the table. Polyhedron Number of Faces Number of Edges Number of Vertices Polyhedron 1 6 12 10 Polyhedron 2 10 16 8 Polyhedron 3 11 17 8 Polyhedron 4 9 15 8 Which polyhedron has incorrect entries for its number of faces, edges, or vertices? Polyhedron 3 Polyhedron 4 Polyhedron 1 Polyhedron 2

OpenStudy (anonymous):

@magbak @Jamierox4ev3r

OpenStudy (ivettef365):

formula F + V - E = 2

OpenStudy (ivettef365):

try to see which one is not true, I got the first one F = 6 E = 12 V = 10 6 + 10 - 12 = 2 16 - 12 = 2 4 = 2 NOT TRUE

OpenStudy (anonymous):

Hm. Thankk youu Ivette(:

OpenStudy (ivettef365):

your welcome, hope you understand :)

OpenStudy (ivettef365):

you can try the others